Respected sir,
[1] The percent profit made when an article is sold for Rs.78 is twice as when it is sold for Rs. 69.The cost price of the article is ?
plz solve my problem...

Answer by stanbon(75887)

The percent profit made when an article is sold for Rs.78 is twice as when it is sold for Rs. 69.The cost price of the article is ?
----------------
Profit = Revenue - Cost
---------
Equations:
2p = 78 - C
p = 69 - C
-----------------
p = 9
----
Substitute into p = 69 - C to get:
9 = 69 - C
Cost = 60Rs
==================
Cheers,
Stan H.